### v1

# Solving the Fokker-Planck equation of discretized Dean-Kawasaki models with functional hierarchical tensor

Abstract
We introduce a novel numerical scheme for solving the Fokker-Planck equation of discretized Dean-Kawasaki models with a functional tensor network ansatz. The Dean-Kawasaki model describes density fluctuations of interacting particle systems, and it is a highly singular stochastic partial differential equation. By performing a finite-volume discretization of the Dean-Kawasaki model, we derive a stochastic differential equation (SDE). To fully characterize the discretized Dean-Kawasaki model, we solve the associated Fokker-Planck equation of the SDE dynamics. In particular, we use a particle-based approach whereby the solution to the Fokker-Planck equation is obtained by performing a series of density estimation tasks from the simulated trajectories, and we use a functional hierarchical tensor model to represent the density. To address the challenge that the sample trajectories are supported on a simplex, we apply a coordinate transformation from the simplex to a Euclidean space by logarithmic parameterization, after which we apply a sketching-based density estimation procedure on the transformed variables. Our approach is general and can be applied to general density estimation tasks over a simplex. We apply the proposed method successfully to the 1D and 2D Dean-Kawasaki models. Moreover, we show that the proposed approach is highly accurate in the presence of external potential and particle interaction.
keywords:
High-dimensional density estimation; Functional tensor network; Dean-Kawasaki model.
1 Introduction
The motivation of this paper is to study the Dean-Kawasaki model [10, 25] with periodic boundary conditions:
(1) $\partial_{t}\pi(x,t)=\frac{1}{\beta}\Delta\pi+\mathrm{div}\left(\sqrt{\frac{2\pi}{\beta N}}\eta+\pi\nabla V_{1}+\pi\nabla V_{2}*\pi\right),\quad\pi(0,x)=\pi_{0}(x),$
where $x\in[0,1]^{n}$ with $n=1,2$, $\eta(x,t)$ is a $\mathbb{R}^{n}$-valued space-time white noise, $N$ controls the strength of the white noise term, $\beta$ is the inverse temperature controlling the diffusion coefficient, $V_{1}(\cdot)$ is an external potential term, and $V_{2}(\cdot)$ is a pairwise potential term modeling particle interactions. Equation 1 is a stochastic partial differential equation (SPDE).
In this work, we study the Dean-Kawasaki model after spatial discretization by dividing the domain into $d$ grid cells. For example, in 1D, the discretization leads to a $d$-dimensional vector $\Pi=(\Pi_{1},\ldots,\Pi_{d})$. Under a discretization scheme such as the finite-volume scheme to be introduced in Section 2, the SPDE in Equation 1 discretizes to a stochastic differential equation (SDE):
(2) $d\Pi(t)=\bm{\mu}(\Pi(t))dt+\bm{\sigma}(\Pi(t))dB_{t}.$
Equation 2 has an associated Fokker-Planck equation.
(3) $\partial_{t}P(\pi,t)=-\mathrm{div}(\bm{\mu}(\pi)P)+\frac{1}{2}\nabla_{\pi}^{2}:(\bm{\Sigma}(\pi)P),\quad\pi\in\Delta,t\in[0,T],$
where $\bm{\Sigma}(\pi)=\bm{\sigma}(\pi)\bm{\sigma}^{\top}(\pi)$ and $\nabla_{\pi}^{2}:(\bm{\Sigma}(\pi)P)=\sum_{ij}\partial_{\pi_{i}}\partial_{\pi_{j}}\left(P(\pi,t)\left(\bm{\Sigma}(\pi)\right)_{ij}\right)$, and $\Delta$ is a simplex which we define later. In particular, a solution $P(\pi,t)$ to Equation 3 is the probability density for $\Pi(t)$ in Equation 2.
In this work, our goal is to solve the high-dimensional partial differential equation (PDE) in Equation 3 with a functional hierarchical tensor ansatz, which has recently shown remarkable success in representing high-dimensional density functions [53, 54]. Due to the difficulty involved in performing time integration on the tensor manifold, this work takes a simple particle-based approach. First, we simulate the SDE in Equation 2 to obtain sample trajectories. Subsequently, we use the functional hierarchical tensor to perform a sketching-based density estimation on the simulated samples. One sees that the approximated density for samples of $\Pi(t)$ exactly corresponds to the solution $P(\cdot,t)$ for Equation 3. Thus, if one is interested in the solution of the Fokker-Planck equation at $t\in[0,T]$, it suffices to perform density estimation tasks based on the samples of $\Pi(t)$. As the density estimation task can be done independently for each time $t$, a resultant appealing property is that there is no additional error accumulation in the proposed particle-based approach.
After the discretization, the distribution of $\Pi(t)$ forms a lattice model in 1D or 2D. After discretization, each SDE trajectory $\Pi(t)$ is supported on a $(d-1)$-dimensional simplex defined as follows:
$\Delta=\{\pi=(\pi_{1},\ldots,\pi_{d})\mid\sum_{i=1}^{d}\pi_{i}=d,0\leq\pi_{i}\leq 1\text{ for }i=1,\ldots,d\},$
which brings a unique challenge to the density estimation task due to the simplex constraint. Moreover, the lattice models have strong site-wise interaction, which makes the model more challenging. To address the simplex constraint, we use a logarithmic transformation to transform the distribution from the $\pi$ variable to a log density $s$. To further improve numerical performance, we use a wavelet transformation and use the wavelet coefficient $c$ as the coordinate. The transformation $\pi\to s\to c$ allows one to have a simplified model for density estimation and improves the efficiency of using the proposed ansatz. We use $P(c,t)$ to encode the pushforward of $P(\pi,t)$ from the transformation $\pi\to c$. The transformation transforms each sample in the $\pi$ coordinate to the $c$ coordinate. As we directly perform density estimation in the $c$ coordinate, there is no need to encode the Jacobian term from the coordinate transformation. In particular, we use the functional hierarchical tensor over a wavelet basis (FHT-W) [54] to perform the density estimation of $P(c,t)$. Our numerical experiments show highly accurate results for several test cases in 1D and 2D Dean-Kawasaki models.
[…]
1.1 Background in the Dean-Kawasaki model
Interacting particle perspective
The Dean-Kawasaki derivation in [10] shows that Equation 1 is associated with the following $N$-particle stochastic differential equation on $X_{1},\ldots,X_{N}$: For $i=1,\ldots,N$, one writes
(4) $dX_{i}(t)=-\left(\nabla V_{1}(X_{i}(t))+\sum_{j=1}^{N}\nabla V_{2}(X_{i}(t)-X_{j}(t))\right)dt+\sqrt{2\beta^{-1}}dB_{i}(t),$
where $B_{1}(t),\ldots,B_{N}(t)$ are $N$ independent Brownian motions in $\mathbb{R}^{n}$. In particular, one sees that Equation 4 is an interacting particle equation with independent diffusion terms. Generalized versions of the Dean-Kawasaki model have found numerous applications in agent-based modeling [19], continuum modeling [32], and they are used to describe bacterial dynamics [55], reaction-diffusion dynamics [27], and nucleation dynamics [35].
Gradient flow perspective
We give some intuitions on how the discretization is derived through a gradient flow characterization of the Dean-Kawasaki model. The diffusion term and the $V_{1},V_{2}$ term in the Dean-Kawasaki model correspond to the energy
[…]
and the Dean-Kawasaki model can be written as
(6) $\partial_{t}\pi(x,t)=\mathrm{div}\left(\pi\nabla\frac{\delta\mathcal{E}(\pi)}{\delta\pi}\right)+\mathrm{div}\left(\sqrt{\frac{2\pi}{\beta N}}\eta\right),$
where $\frac{\delta\mathcal{E}(\pi)}{\delta\pi}$ is the Fréchet derivative. One can see that Equation 6 indicates that the Dean-Kawasaki model is a stochastic Wasserstein gradient flow. In particular, the $\bm{\mu}$ term in Equation 2 is obtained by taking a cell average of the deterministic flux term $\pi\nabla\frac{\delta\mathcal{E}(\pi)}{\delta\pi}$, and the $\bm{\sigma}$ term in Equation 2 is obtained by taking a cell average of the stochastic flux term $\sqrt{\pi}\eta$. This gradient flow characterization is essential to deriving the discrete Dean-Kawasaki model, and this approach is also taken by [42].
1.2 Related work
Density estimation in Lattice models
There exists a significant body of research on density estimation subroutines applicable to lattice models. In the continuous case, notable neural network models include restricted Boltzmann machines (RBMs) [20, 43], energy-based models (EBMs) [21, 31, 16, 36], normalizing flows [49, 48, 40, 37], variational auto-encoders (VAEs) [12, 28], diffusion and flow-based models [44, 56, 46, 47, 45, 22, 2, 34, 33, 1]. Tensor network models have recently emerged as a suitable density estimation model over the continuous lattice models [5, 53, 52, 54]. For discrete density estimation on lattice models, tensor network models provide a direct representation of high-dimensional discrete distributions, and notable examples include [17, 6, 14, 4, 15, 51, 39, 26, 38].
The Dean-Kawasaki model
The Dean-Kawasaki model has long been of research interest in non-equilibrium statistical physics due to its direct modeling of microscopic interacting particle dynamics. The Dean-Kawasaki model is also an example of fluctuating hydrodynamics [18, 13, 3, 11, 23]. It is noteworthy that the Dean-Kawasaki model does not admit martingale solutions for generic choices of parameters and initial condition, which was first proven for a special potential-free case in [30] and in [29] for the general case. The well-posedness of the Dean-Kawasaki model does not directly relate to our work, as the spatial discretization is sufficient to provide an implicit regularization to the Dean-Kawasaki model. Further regularization has been proposed to ensure the stability of the Dean-Kawasaki model [9, 8, 7]. As our approach is suitable for density estimation over a general simplex, our methodology extends to the regularized Dean-Kawasaki model.
2 Discretization of the Dean-Kawasaki model
We first describe the approach of performing spatial discretization of the Dean-Kawasaki model to form a finite-dimensional SDE. For simplicity, we consider an $n$D case with $n=1,2$, and the domain is an $n$-dimensional cube $\Omega=[0,1]^{n}$ prescribed with periodic boundary condition in each physical dimension. The finite-volume treatment generalizes to high-dimensional cases, which we omit for simplicity. Our simple approach is motivated by the finite-volume approaches adopted in [27, 42]. In general, any reasonable finite-volume discretization can be used, and they only result in a modification of the associated Fokker-Planck equation in Equation 3. Therefore, our subsequent tensor network density estimation approach would accommodate other discretization procedures.
We give an outline for this section. Section 2.1 provides a discretization of the Dean-Kawasaki model in 1D. In Section 2.2, we carry out the derivation in 2D, which is more notationally involved and can be skipped for the first read.
2.1 1D discretization
We give the derivation for the finite-volume discretization to simulate the 1D Dean-Kawasaki model in Equation 1. The goal is to derive an SDE as shown in Equation 2, which would then lead to the Fokker-Planck equation in Equation 3.
We take $h=\frac{1}{m}$ to be the spatial grid mesh size and take $x_{j}=(j-\frac{1}{2})h$ for $j=1,\ldots,m$ to be the cell center points. We take $x_{j+\frac{1}{2}}=jh$ to be the cell boundary points. The domain is subdivided into grid cells $C_{j}=[x_{j-\frac{1}{2}},x_{j+\frac{1}{2}}]=[(j-1)h,jh]$. The finite-volume approach takes a cell average of $\pi$ over each $C_{j}$. Formally, we write
[…]
and $\Pi(t)$ is a $d$-dimensional discretization of $\pi$ with $d=m$.
By the discussion in Section 1.1, we see that the Dean-Kawasaki model is a stochastic gradient flow driven by a deterministic and stochastic flux term. For $\mathcal{E}$ as defined in Equation 5, one has the two flux term written as follows
$F_{d}(x,t)=\pi\nabla\frac{\delta\mathcal{E}(\pi)}{\delta\pi},\quad F_{s}(x,t)=\sqrt{\frac{2\pi}{\beta N}}\eta,$
where $\eta$ is a spatiotemporal white noise, $F_{d}$ is the deterministic flux, and $F_{s}$ is the stochastic flux.
One can then derive the discretized Dean-Kawasaki model by integrating Equation 6. By the divergence theorem, one has
$d\Pi_{j}(t)=\frac{1}{h}\left(F_{d,j+\frac{1}{2}}-F_{d,j-\frac{1}{2}}\right)+\frac{1}{h}\left(F_{s,j+\frac{1}{2}}-F_{s,j-\frac{1}{2}}\right),$
where $F_{d,j+\frac{1}{2}},F_{s,j+\frac{1}{2}}$ are the evaluation of $F_{d}$ and $F_{s}$ at the boundary $x_{j+\frac{1}{2}}$. In particular, the approximation for $F_{d,j+\frac{1}{2}},F_{d,j-\frac{1}{2}}$ forms the $\bm{\mu}$ term in Equation 2, and the approximation for $F_{s,j+\frac{1}{2}},F_{s,j-\frac{1}{2}}$ forms the $\bm{\sigma}$ term in Equation 2.
To further simplify the equation, we use $F_{V,j+\frac{1}{2}}$ to denote the flux at the boundary $x_{j+\frac{1}{2}}$ corresponding to the $V_{1},V_{2}$ terms in $\mathcal{E}$. For the diffusion term corresponding to the entropy term in $\mathcal{E}$, we use a simple central difference scheme, leading to an SDE
(7) $d\Pi_{j}(t)$ $=\frac{1}{\beta h^{2}}\left(\Pi_{j+1}(t)+\Pi_{j-1}(t)-2\Pi_{j}(t)\right)dt$
$+\frac{1}{h}\left(F_{V,j+\frac{1}{2}}-F_{V,j-\frac{1}{2}}\right)+\frac{1}{h}\left(F_{s,j+\frac{1}{2}}-F_{s,j-\frac{1}{2}}\right).$
First, we describe how one approximates $F_{V,j+\frac{1}{2}}$. We use the Roe scheme [41]. Define $F_{V,j}$ to be the flux term through $x_{j}$ corresponding to the $V_{1},V_{2}$ term in $\mathcal{E}$. The discretization takes the form
$F_{V,j}=\Pi_{j}(t)V_{1}^{\prime}(x_{j})dt+\frac{1}{d}\sum_{i}V_{2}^{\prime}(x_{j}-x_{i})\Pi_{i}(t)\Pi_{j}(t)dt.$
Then, for the $V_{1},V_{2}$ term, the corresponding flux term is defined by
$F_{V,j+\frac{1}{2}}=\begin{cases}F_{V,j}&A_{V,j+\frac{1}{2}}<0,\\ F_{V,j+1}&A_{V,j+\frac{1}{2}}\geq 0,\end{cases}$
where
$A_{V,j+\frac{1}{2}}=\frac{F_{V,j+1}-F_{V,j}}{\Pi_{j+1}(t)-\Pi_{j}(t)}.$
For the stochastic flux term $F_{s,j+\frac{1}{2}}$, we take the approach described in [27]. Let $B_{j+\frac{1}{2}}(t)$ be an independent Wiener process for $j=0,\ldots,d-1$. The discretization takes the form
$F_{s,j+\frac{1}{2}}=\sqrt{\frac{2\tilde{\Pi}_{j+\frac{1}{2}}(t)}{h\beta N}}dB_{j+\frac{1}{2}}(t),$
where $\tilde{\Pi}_{j+\frac{1}{2}}$ is the spatial average of $\pi(x,t)$ around $x=x_{j+\frac{1}{2}}$. In particular, $\tilde{\Pi}_{j+\frac{1}{2}}$ is a modified arithmetic mean defined by
$\tilde{\Pi}_{j+\frac{1}{2}}(t)=\frac{\Pi_{j}(t)+\Pi_{j+1}(t)}{2}H\left(hN\Pi_{j}(t)\right)H\left(hN\Pi_{j+1}(t)\right),$
where $H$ is a smoothed Heaviside function defined by the following form:
$H(x)=\begin{cases}0&x\leq 0,\\ x&0\leq x\leq 1,\\ 1&x\geq 1.\end{cases}$
The definition of $F_{s,j+\frac{1}{2}}$ is the approach taken in [27], and a particular advantage of this construction is numerical stability. When $V_{1}=V_{2}=0$, it has been shown in [27] that the resulting continuous-time SDE in Equation 7 preserves the positivity of $\Pi_{j}(t)<0$ for any $j,t$.
The aforementioned definition of $F_{V,j+\frac{1}{2}}$ and $F_{s,j+\frac{1}{2}}$ fully characterizes the SDE in Equation 7. Overall, one sees that the finite-volume discretization allows one to write the SPDE in Equation 1 as a tractable SDE in Equation 7. To simulate the SDE, we further perform a temporal discretization. Let $T$ be the terminal time. We take $\delta t$ to be a fixed time grid mesh size and $K=\lceil T/\delta t\rceil+1$. We define $t_{k}=(k-1)\delta t$ for $k=1,\ldots,K$. We take $\Pi^{k}_{j}$ to model $\Pi_{j}(t_{k})$.
[…]
(8) $\Pi_{j}^{k+1}=$ $\Pi_{j}^{k}+\frac{\delta t}{\beta h^{2}}\left(\Pi^{k+1}_{j+1}+\Pi^{k+1}_{j-1}-2\Pi^{k+1}_{j}\right)$
$+$ $\frac{\delta t}{h}\left(F^{k}_{V,j+\frac{1}{2}}-F^{k}_{V,j-\frac{1}{2}}\right)+\frac{\sqrt{\delta t}}{h}\left(F^{k}_{s,j+\frac{1}{2}}-F^{k}_{s,j-\frac{1}{2}}\right),$
where we use an implicit scheme for the diffusion term $\Delta\pi$ to ensure stability under larger $\delta t$.
Thus, it remains to fully specify $F^{k}_{V,j+\frac{1}{2}}$ and $F^{k}_{s,j+\frac{1}{2}}$. We first specify $F^{k}_{V,j+\frac{1}{2}}$. We use $F_{V,j}^{k}$ to represent $F_{V,j}$ at time $t_{k}$. We write
$F_{V,j}=\Pi_{j}^{k}V_{1}^{\prime}(x_{j})+\frac{1}{d}\Pi_{j}^{k}\sum_{i}V_{2}^{\prime}(x_{j}-x_{i})\Pi_{i}^{k}.$
Then, we define $F^{k}_{V,j+\frac{1}{2}}$ by the following:
$F^{k}_{V,j+\frac{1}{2}}=\begin{cases}F^{k}_{V,j}&A^{k}_{V,j+\frac{1}{2}}<0,\\ F^{k}_{V,j+1}&A^{k}_{V,j+\frac{1}{2}}\geq 0,\end{cases}\quad A^{k}_{V,j+\frac{1}{2}}=\frac{F_{V,j+1}^{k}-F_{V,j}^{k}}{\Pi_{j+1}^{k}-\Pi_{j}^{k}}.$
We then specify $F^{k}_{s,j+\frac{1}{2}}$. Let $W^{k}_{j+\frac{1}{2}}\sim N(0,1)$ be i.i.d. standard normal distribution for $k=1,\ldots,K,j=0,\ldots,m-1$. We write
$F_{s,j+\frac{1}{2}}^{k}=\sqrt{\frac{2\tilde{\Pi}^{k}_{j+\frac{1}{2}}}{h\beta}}W_{j+\frac{1}{2}}^{k},\quad\tilde{\Pi}^{k}_{j+\frac{1}{2}}=\frac{\Pi_{j}^{k}+\Pi_{j+1}^{k}}{2}H\left(hN\Pi_{j}^{k}\right)H\left(hN\Pi_{j+1}^{k}\right).$
Thus, simulation with respect to the SDE in Equation 7 allows one to simulate the discretized Dean-Kawasaki model. For any $t=t_{k}$, the simulation results of $(\Pi^{k}_{j})_{j=1}^{d}$ forms an approximation to the SPDE solution $\pi(\cdot,t_{k})$. For practical purposes, we perform a clamping on the Gaussian variable $W_{j+\frac{1}{2}}^{k}$. We replace $W_{j+\frac{1}{2}}^{k}$ with $\max(-5,\min(W_{j+\frac{1}{2}}^{k},5))$ to prevent blow-up from sampling an overly large normal random variable. In practice, a more careful treatment with noise clamping can be done for future research.
2.2 2D discretization
A similar procedure applies to 2D by modest modification to the 1D case. For simplicity, we only define the spatial discretization of the SDE. The temporal discretization follows from a simple modification of Equation 8. We take $h=\frac{1}{m}$ and take $x_{j}=(j-\frac{1}{2})h$ for $j=1,\ldots,m$. We take $x_{j+\frac{1}{2}}=jh$. The domain is subdivided into grid cells $C_{(i,j)}=[x_{i-\frac{1}{2}},x_{i+\frac{1}{2}}]\times[x_{j-\frac{1}{2}},x_{j+\frac{1}{2}}]=[(i-1)h,ih]\times[(j-1)h,jh]$. The finite-volume approach takes a cell average of $\pi$ over each $C_{i,j}$. Formally, we write
[…]
and $\Pi(t)$ is a $d$-dimensional discretization of $\pi$ with $d=m^{2}$.
In particular, the gradient flow perspective in Section 2.1 carries through. One writes
$(F_{d,1}(x,t),F_{d,2}(x,t))=\pi\nabla\frac{\delta\mathcal{E}(\pi)}{\delta\pi},\quad(F_{s,1}(x,t),F_{s,2}(x,t))=\sqrt{\frac{2\pi}{\beta N}}\eta,$
and the divergence theorem leads to
$d\Pi_{(i,j)}(t)=$ $\frac{1}{h}\left(F_{d,1,(i+\frac{1}{2},j+\frac{1}{2})}-F_{d,1,(i-\frac{1}{2},j+\frac{1}{2})}\right)+\frac{1}{h}\left(F_{s,1,(i+\frac{1}{2},j+\frac{1}{2})}-F_{s,1,(i-\frac{1}{2},j+\frac{1}{2})}\right)$
$+$ $\frac{1}{h}\left(F_{d,2,(i+\frac{1}{2},j+\frac{1}{2})}-F_{d,2,(i+\frac{1}{2},j-\frac{1}{2})}\right)+\frac{1}{h}\left(F_{s,2,(i+\frac{1}{2},j+\frac{1}{2})}-F_{s,2,(i+\frac{1}{2},j-\frac{1}{2})}\right),$
where $F_{d,1,(i+\frac{1}{2},j+\frac{1}{2})},F_{s,1,(i+\frac{1}{2},j+\frac{1}{2})}$ are respectively the average flux of $F_{d,1},F_{s,1}$ into the right boundary of $C_{(i,j)}$, and $F_{d,2,(i+\frac{1}{2},j+\frac{1}{2})},F_{s,2,(i+\frac{1}{2},j+\frac{1}{2})}$ are respectively the average flux of $F_{d,2},F_{s,2}$ into the upper boundary of $C_{(i,j)}$. To formally define $F_{d,1,(i+\frac{1}{2},j+\frac{1}{2})}$, for example, one writes
[…]
and the rest of the formal definition follows likewise.
Similar to the 1D case, we separate the deterministic flux term by using a central difference scheme to approximate the diffusion term. For $l=1,2$, we use $F_{V,l,(i+\frac{1}{2},j+\frac{1}{2})}$ to represent the part of $F_{d,l,(i+\frac{1}{2},j+\frac{1}{2})}$ corresponding to the $V_{1},V_{2}$ term. Then, in this case, the SDE writes
(9) $d\Pi_{(i,j)}(t)$
$=$ $\frac{\delta t}{\beta h^{2}}\left(\Pi_{(i+1,j)}(t)+\Pi_{(i-1,j)}(t)+\Pi_{(i,j+1)}(t)+\Pi_{(i,j-1)}(t)-4\Pi_{(i,j)}(t)\right)$
$+$ $\frac{1}{h}\left(F_{V,1,(i+\frac{1}{2},j+\frac{1}{2})}-F_{V,1,(i-\frac{1}{2},j+\frac{1}{2})}\right)+\frac{1}{h}\left(F_{s,1,(i+\frac{1}{2},j+\frac{1}{2})}-F_{s,1,(i-\frac{1}{2},j+\frac{1}{2})}\right)$
$+$ $\frac{1}{h}\left(F_{V,2,(i+\frac{1}{2},j+\frac{1}{2})}-F_{V,2,(i+\frac{1}{2},j-\frac{1}{2})}\right)+\frac{1}{h}\left(F_{s,2,(i+\frac{1}{2},j+\frac{1}{2})}-F_{s,2,(i+\frac{1}{2},j-\frac{1}{2})}\right).$
First, we describe how one approximates $F_{V,l,(i+\frac{1}{2},j+\frac{1}{2})}$. We use a Roe scheme. Define $F_{V,l,(i,j)}$ to be the flux term through $x_{(i,j)}$ corresponding to to the $V_{1},V_{2}$ term in $\mathcal{E}$. The discretization takes the following form for $l=1,2$:
$F_{V,l,(i,j)}=$ $\Pi_{(i,j)}(t)\partial_{x_{l}}V_{1}(x_{(i,j)})+\frac{1}{d}\sum_{i^{\prime},j^{\prime}}\partial_{x_{l}}V_{2}(x_{(i,j)}-x_{(i^{\prime},j^{\prime})})\Pi_{(i^{\prime},j^{\prime})}(t)\Pi_{(i,j)}(t).$
Then, for $l=1,2$, the numerical flux is defined by
$F_{V,l,(i+\frac{1}{2},j+\frac{1}{2})}=\begin{cases}F_{V,l,(i,j)}&A_{V,l,(i+\frac{1}{2},j+\frac{1}{2})}<0,\\ F_{V,l,(i+1,j)}&A_{V,l,(i+\frac{1}{2},j+\frac{1}{2})}\geq 0,l=1,\\ F_{V,l,(i,j+1)}&A_{V,l,(i+\frac{1}{2},j+\frac{1}{2})}\geq 0,l=2,\end{cases}$
where
$A_{V,1,(i+\frac{1}{2},j+\frac{1}{2})}=\frac{F_{V,1,(i+1,j)}-F_{V,1,(i,j)}}{\Pi_{(i+1,j)}(t)-\Pi_{(i,j)}(t)},\quad A^{k}_{V,2,(i+\frac{1}{2},j+\frac{1}{2})}=\frac{F_{V,2,(i,j+1)}-F_{V,2,(i,j)}}{\Pi_{(i,j+1)}(t)-\Pi_{(i,j)}(t)}.$
For the stochastic flux term $F_{s,l,(i+\frac{1}{2},j+\frac{1}{2})}$, we likewise use the scheme in [27]. Let $B_{l,(i+\frac{1}{2},j+\frac{1}{2})}(t)$ be independent Wiener process for $i,j=0,\ldots,m-1$ and $l=1,2$. Then, the discretization takes the form
$F_{s,l,(i+\frac{1}{2},j+\frac{1}{2})}=\sqrt{\frac{2\tilde{\Pi}_{l,(i+\frac{1}{2},j+\frac{1}{2})}(t)}{h^{2}\beta N}}dB_{l,(i+\frac{1}{2},j+\frac{1}{2})}(t),$
where $\tilde{\Pi}_{1,(i+\frac{1}{2},j+\frac{1}{2})}(t)$ and $\tilde{\Pi}_{2,(i+\frac{1}{2},j+\frac{1}{2})}(t)$ are both spatial approximation of $\pi$ around $(x_{i+\frac{1}{2}},x_{j+\frac{1}{2}})$. Following the construction in [27], we take
$\tilde{\Pi}_{1,(i+\frac{1}{2},j+\frac{1}{2})}(t)=\frac{\Pi_{(i,j)}^{k}+\Pi_{(i+1,j)}^{k}}{2}H\left(h\Pi_{(i,j)}^{k}\right)H\left(h\Pi_{(i+1,j)}^{k}\right),$
and
$\tilde{\Pi}_{2,(i+\frac{1}{2},j+\frac{1}{2})}(t)=\frac{\Pi_{(i,j)}^{k}+\Pi_{(i,j+1)}^{k}}{2}H\left(h\Pi_{(i,j)}^{k}\right)H\left(h\Pi_{(i,j+1)}^{k}\right),$
where $H(\cdot)$ is the smoothed heaviside function as defined in the 1D case.
Simulation with respect to the SDE in Equation 9 allows one to simulate the discretized Dean-Kawasaki model. For any $t=t_{k}$, the simulation results of $(\Pi^{k}_{(i,j)})_{i,j=1}^{m}$ forms an approximation to the SPDE solution $\pi(\cdot,t_{k})$. For the 2D case, we also use clamping to ensure stability.
3 Main formulation
This section details the approach to perform density estimation over a $(d-1)$-dimensional simplex $\Delta$ with a functional hierarchical tensor. Our goal is to solve the Fokker-Planck equation in Equation 3 with density estimation. Thus, the input to this subroutine is a collection of samples on $\Delta$ at time $t$, and the output is a normalized density $P(\pi,t)$ in a functional tensor network ansatz.
Explicitly, by simulating the SDE in Equation 7 and Equation 9, we obtain a collection of samples. For the case of 1D, we obtain a collection of $B$ samples $\{\Pi^{(b)}(t)=(\Pi^{(b)}_{j}(t))_{j=1,\ldots,m}\}_{b=1,\ldots,B}$. Likewise, in the case of 2D, we obtain a collection of $B$ samples $\{\Pi^{(b)}(t)=(\Pi^{(b)}_{(i,j)}(t))_{i,j=1,\ldots,m}\}_{\,b=1,\ldots,B}$. The samples at time $t$ forms a empirical distribution approximation to the Fokker-Planck equation solution $P(\pi,t)\approx\frac{1}{N}\sum_{b=1}^{B}\delta(\pi-\Pi^{(b)}(t))$. In other words, the discretized Dean-Kawasaki model introduced in Section 2 describes a specific probability distribution over $\Delta$, and performing the density estimation over the simulated samples can give a particle-based approximation to the associated Fokker-Planck equation in Equation 3.
We go through the outline of this section by expanding on how we address a few challenges in the aforementioned density estimation approach. First, conventional tensor network density estimation only applies to Euclidean spaces, and the simplex constraint in Equation 10 for the Fokker-Planck equation is difficult for tensor networks. To address the issue, we use an entry-wise logarithmic transformation followed by an entry-wise offset to center the distribution. Overall, this coordinate transformation maps $\Delta$ to $\mathbb{R}^{d}$ and effectively eliminates the simplex constraints. This coordinate transformation is covered in the first part of Section 3.1.
Following the logarithmic coordinate transformation, one obtains a lattice model, and a challenge is that the lattice model exhibits strong site-wise coupling for practically interesting models. To efficiently address the strong site-wise coupling, we perform a wavelet transformation that separates the model into different length scales. The coordinate transformation to the wavelet coefficient is covered in the second part of Section 3.1. After the transformation to the wavelet coefficient space, we use the functional hierarchical tensor over a wavelet basis (FHT-W) ansatz for density estimation. The FHT-W model is covered in Section 3.2.
Lastly, the transformation into the wavelet coefficient leads to challenges in obtaining accurate observable estimation. Essentially, when one has obtained the density function in the form of an FHT-W ansatz, one also needs to form the observable function in the FHT-W ansatz to perform accurate observable estimation. We propose to use interpolation for compressing the observable function in the FHT-W ansatz. The interpolation procedure is covered in Section 3.3, which is technical and can be skipped for a first reading.
Notation
For notational compactness, we introduce several shorthand notations for simple derivation. For $n\in\mathbb{N}$, let $[n]:=\{1,\ldots,n\}$. For an index set $S\subset[d]$, we let $x_{S}$ stand for the subvector with entries from index set $S$.
3.1 Mapping $\Delta$ to Euclidean space
We shall describe the an invertible transformation of $\Delta$ to $\mathbb{R}^{d-1}$ through a coordinate transformation from $\pi$ to $c$. This procedure is done by two coordinate transformations $\pi\to s$ followed by $s\to c$, which we shall cover sequentially.
Transform to the log space
The $(d-1)$-dimensional simplex $\Delta$ is defined as follows:
(10) $\Delta=\{\pi=(\pi_{1},\ldots,\pi_{d})\mid\sum_{i=1}^{d}\pi_{i}=d,0\leq\pi_{i}\leq 1\text{ for }i=1,\ldots,d\},$
and we see that $\Delta\subset\mathbb{R}^{d}$ is a polytope. It is less suitable to directly perform density estimation of $\Delta$ in the $\pi$ coordinate. Instead, we propose to apply a nonlinear transformation of $\Delta$ as a subset of $\mathbb{R}^{d}$. We define the map $s\colon\Delta\to\mathbb{R}^{d}$ by the following formula:
(11) $s(\pi_{1},\ldots,\pi_{d})=(\log(\pi_{1})-Y,\ldots,\log(\pi_{d})-Y),\quad Y=\frac{1}{d}\sum_{j=1}^{d}\log(\pi_{j}).$
We explain how Equation 11 is designed. The formula for $s$ consists of two steps: (a) perform entry-wise logarithm to get $\log(\pi)$, (b) subtract by the mean to get $\log(\pi)-Y$. We write $s(\pi)=s=(s_{1},\ldots,s_{d})$. This nonlinear transformation transforms $\Delta$ from the $\pi$ coordinate to the $s$ coordinate. One can see that $s(\pi)$ is the log density of $\pi$. The map $\pi\to(s_{1},\ldots,s_{d-1})$ is an embedding of $\Delta$ into an open subset of $\mathbb{R}^{d-1}$.
We see that the $\pi$ to $s$ transformation already allows a conventional density estimation subroutine to be applicable. Let $P(\pi,t)$ denote the target distribution over $\Delta$ under the $\pi$ coordinate. With the $s$ transformation, we can instead consider $P(s,t)$, which is the target distribution under the $s$ coordinate. Formally, $P(s,t)$ is the pushforward of $P(\pi,t)$ under $s(\cdot)$. Moreover, one is given samples to $P(s,t)$ by transforming the samples of $P(\pi,t)$ through $s(\cdot)$. One can directly apply the FHT ansatz in [53] to perform density estimation from samples of $P(s,t)$. The issue that $P(s,t)$ is a singular measure can be easily addressed by performing the density estimation of the marginal distribution of $P(s,t)$ over the first $d-1$ entries.
Lastly, for $s=(s_{1},\ldots,s_{d})$, one can define an inverse map $s^{-1}$ by
$s^{-1}(s_{1},\ldots,s_{d})=\left(\exp(s_{1}-W),\ldots,\exp(s_{d}-W)\right),$
where $W=-\log(d)+\mathrm{log}\left(\sum_{j=1}^{d}\exp(s_{j})\right)$.
Transform to the wavelet space
We perform a further coordinate transformation through wavelet transformation. In our discretization for the Dean-Kawasaki model, the distribution $P(\pi,t)$ is a lattice model in 1D and 2D. With the pushforward into the log space, $P(s,t)$ is also a lattice model. In practice, $P(s,t)$ has a strong site-wise interaction, and a wavelet transformation $s\to c$ essentially serves to reduce the numerical rank in the new target distribution $P(c,t)$, and indeed the wavelet transformation is shown in [54] to be effective for several lattice models with strong site-wise correlation. Moreover, the constraint that entries of $s$ sum to zero can be elegantly addressed by the wavelet transformation.
We first go through the iterative wavelet coarsening procedure in the 1D case. For simplicity, we assume that $d=2^{L}$ so that $\pi=(\pi_{1},\ldots,\pi_{2^{L}})$. In the first step of the coarse-graining process, the wavelet filter transforms the $s$ variable into two $2^{L-1}$-dimensional variables $y_{L-1}=(y_{j,L-1})_{j\in[2^{L-1}]}$ and $c_{L-1}=(c_{j,L-1})_{j\in[2^{L-1}]}$, which are defined by the following equation
$y_{j,L-1}=\frac{s_{2j-1}+s_{2j}}{\sqrt{2}},\quad c_{j,L-1}=\frac{s_{2j-1}-s_{2j}}{\sqrt{2}}.$
The variable $y_{L-1}$ is the scaling coefficient of $s$ at length scale $2^{-(L-1)}$, and $c_{L-1}$ is the detail coefficient of $s$ at length scale $2^{-(L-1)}$.
Repeating the same approach, for $l=L-1,\ldots,1$, one performs wavelet transformation on $y_{l}$ to obtain $2^{l-1}$-dimensional variables $y_{l-1}$ and $c_{l-1}$ given by the following equation
$y_{j,l-1}=\frac{y_{2j-1,l}+y_{2j,l}}{\sqrt{2}},\quad c_{j,l-1}=\frac{y_{2j-1,l}-y_{2j,l}}{\sqrt{2}}.$
Finally, we note that $y_{1}$ is $2$-dimensional the procedure ends at $y_{1}\to(y_{0},c_{0})$. As $\sum_{j=1}^{2^{L}}s_{j}=0$ due to Equation 11, we have $y_{0}=0$. Therefore, we have fully described the procedure to transform $s=(s_{1},\ldots,s_{d})$ into $(c_{k,l})_{k,l}$. In particular, the transformation from $\pi$ to $s=(s_{1},\ldots,s_{d})$ to $(c_{k,l})_{l=0,\ldots,L,\,k=1,\ldots,2^{l}}$ is an invertible embedding from $\Delta$ to an subset of $\mathbb{R}^{d-1}$.
By the same procedure, one can perform the iterative wavelet transformation on the $s$ variable in the 2D case. One can perform a series of iterative 2D wavelet transformation to the $s$ variable $s=(s_{(i,j)})_{i,j=1}^{m}$ with $m=2^{L/2}$. By the procedure in [54], one likewise transforms the $s$ variable to $(c_{k,l})_{l=0,\ldots,L,\,k=1,\ldots,2^{l}}$. The map from $\pi$ to $c$ is also invertible in this case.
Thus, for time $t\in[0,T]$, one is given a collection of the $B$ samples $\{\Pi^{(b)}(t)=(\Pi^{(b)}_{j}(t))_{j=1,\ldots,m}\}_{b=1,\ldots,B}$ in 1D or $\{\Pi^{(b)}(t)=(\Pi^{(b)}_{(i,j)}(t))_{i,j=1,\ldots,m}\}_{b=1,\ldots,B}$ in 2D, the $s$ transformation composed with the 1D/2D wavelet transformation results in a collection of samples in the $c$ space. We denote the samples in wavelet coefficient as $\{C^{(b)}(t)=(C^{(b)}_{k,l}(t))_{l=0,\ldots,L,\,k=1,\ldots,2^{l}}\}_{b=1,\ldots,B}$. The wavelet transformed samples $\{C^{(b)}(t)\}_{b=1,\ldots,B}$ are the inputs to the subsequent density estimation subroutine. Moreover, the samples form an empirical distribution approximation to $P(c,t)$. We subsequently use a tensor network ansatz to perform density estimation over the samples, which is essentially an approximation of the empirical distribution with a tensor network ansatz.
3.2 Tensor network architecture of $P(s,t)$
After the transformation from the $\pi$ coordinate into the wavelet coordinate $(c_{k,l})_{k,l}$, we use a functional hierarchical tensor over a wavelet basis (FHT-W) ansatz to model the distribution of $P(c,t)$. The $l=0,\ldots,L-1$ index is the resolution scale index, with a larger level $l$ standing for information on a finer scale. The $k=1,\ldots,2^{l}$ index is the location index at level $l$. Given the $(c_{k,l})_{l=0,\ldots,L,\,k=1,\ldots,2^{l}}$ variable in the wavelet coordinate, one can directly use the FHT-W model, illustrated in Figure 1. The FHT-W model is a tree-based functional tensor network based on a tree structure obtained from a modified binary tree.
Overall, the TN architecture for FHT-W preserves the hierarchical structure of $(c_{k,l})_{l=0,\ldots,L,\,k=1,\ldots,2^{l}}$, and internal nodes are inserted to ensure computational efficiency. A description of the FHT-W and the density estimation procedure can be found in [54].
Figure 1: Illustration of the FHT-W ansatz for $L=4$. Each $v_{k,l}$ represents the tensor component $G_{v_{k,l}}$ at the external node $v_{k,l}$, and the $v_{k,l}$ corresponds to the variable $c_{k,l}$ in the wavelet transformation. Each $w_{k,l}$ represents the tensor component $G_{w_{k,l}}$ at the internal node $w_{k,l}$.
To unfamiliar readers, we now cover the essential backgrounds in tree-based functional tensor networks necessary to understand FHT-W. Details can be found in [54].
Notations for tree graphs
A tree graph $T=(V,E)$ is a connected undirected graph without cycles. An undirected edge $\{v,v^{\prime}\}\in E$ is written interchangeably as $(v,v^{\prime})$ or $(v^{\prime},v)$. For any $v\in V$, define $\mathcal{N}(v)$ to be the neighbors of $v$. Moreover, define $\mathcal{E}(v)$ as the set of edges incident to $v$. For an edge $e=(k,w)$, removing the edge $e$ in $E$ results in two connected components $I_{1},I_{2}$ with $I_{1}\cup I_{2}=V$. For any $(w,k)\in E$, we use $k\to w$ to denote the unique connected component among $I_{1},I_{2}$ which contains $k$ as a node.
Tree-based FTN
We introduce the tree-based functional tensor network ansatz to be used for density estimation. Let $T=(V,E)$ be a tree graph and let $V_{\mathrm{ext}}$ be the subset of vertices in $T$ whose associated tensor component admits an external bond. To adhere to the setting of $d$-dimensional functions, let $|V_{\mathrm{ext}}|=d$. The total vertex size $\tilde{d}=|V|$ is equal to $d$ only when the tensor network has no internal nodes. We define the tree-based functional tensor network in Definition 3.1. The tensor network is illustrated in Figure 2
Definition 3.1.
(Tree-based functional tensor network) Suppose one has a tree structure $T=(V,E)$ and the corresponding ranks $\{r_{e}:e\in E\}$. The symbol $V_{\mathrm{ext}}$ is the subset of $V$ that contains external bonds. For simplicity, we apply a labeling of nodes so that $V=[\tilde{d}]$ and $V_{\mathrm{ext}}=[d]$.
The tensor component at $k\in V$ is denoted by $G_{k}$. Let $\mathrm{deg}(k)$ stand for the degree of $k$ in $T$. When $k\in V_{\mathrm{ext}}$, $G_{k}$ is defined as a $(\mathrm{deg}(k)+1)$-tensor of the following shape:
$G_{k}:[n_{k}]\times\prod_{e\in\mathcal{E}(k)}[r_{e}]\rightarrow\mathbb{R}.$
When $k\not\in V_{\mathrm{ext}}$, $G_{k}$ is defined as a $\mathrm{deg}(k)$-tensor of the following shape:
$G_{k}:\prod_{e\in\mathcal{E}(k)}[r_{e}]\rightarrow\mathbb{R}.$
A $d$-tensor $D$ is said to be a tree tensor network defined over the tree $T$ and tensor components $\{G_{k}\}_{k=1}^{\tilde{d}}$ if
(12) $D(i_{1},\ldots,i_{d})=\sum_{\alpha_{E}}\prod_{k=1}^{d}G_{k}\left(i_{k},\alpha_{\mathcal{E}(k)}\right)\prod_{k=d+1}^{\tilde{d}}G_{k}\left(\alpha_{\mathcal{E}(k)}\right).$
Moreover, for $j=1,\ldots,d$, let $\{\psi_{i,j}\}_{i=1}^{n_{j}}$ denote a collection of orthonormal function basis over a single variable for the $j$-th coordinate. A tree-based FTN is the $d$-dimensional function defined by $D$, and it admits the following equation:
(13) $p(x_{1},\ldots,x_{d})=\sum_{\alpha_{E}}\prod_{k=1}^{d}\left(\sum_{i_{k}=1}^{n_{k}}G_{k}\left(i_{k},\alpha_{\mathcal{E}(k)}\right)\psi_{i_{k},k}(x_{k})\right)\prod_{k=d+1}^{\tilde{d}}G_{k}\left(\alpha_{\mathcal{E}(k)}\right).$
Figure 2: (A) A tree structure $T=(V,E)$ with $V=V_{\mathrm{ext}}=\{1,\ldots,10\}$. (B) Tensor Diagram representation of a tree tensor network over $T$.
Tree structure for FHT-W
We first specify the vertex of $T$ in the case of the FHT-W model used for the discretized Dean-Kawasaki model SDE. We remark that the data structure of the $c=(c_{k,l})_{k,l}$ variable is the same for both the 1D and 2D cases. For the vertex set, we define the external nodes to be $V_{\mathrm{ext}}=\{v_{k,l}\}_{l\in[L-1]\cup\{0\},k\in[2^{\max(l,0)}]}$. The total vertex set is $V=V_{\mathrm{ext}}\cup\{w_{k,l}\}_{l\in[L-2]\cup\{0\},k\in[2^{l}]}$. The vertex $v_{k,l}$ corresponds to the variable $c_{k,l}$, and the variables $w_{k,l}$ is an internal node inserted at each level $l=0,\ldots,L-2$.
The edge set of $T$ is $E=E_{1}\cup E_{2}$, where $E_{1}=\{(w_{k,l},v_{k,l})\}_{w_{k,l}\in V}$ and $E_{2}=\{(w_{k,l},v_{2k-1,l+1}),(w_{k,l},v_{2k,l+1})\}_{w_{k,l}\in V}$. The definition of $E$ shows that each variable $c_{2k-1,l+1}$ and $c_{2k,l+1}$ is connected to the variable $c_{k,l}$ through an internal node $w_{k,l}$. This construction for $T$ ensures that $c_{k,l}$ is placed closed to $c_{k^{\prime},l^{\prime}}$ when $(k,l)$ is close to $(k^{\prime},l^{\prime})$. The description fully characterizes the tree $T=(V,E)$. One sees that the description aligns with the illustration in Figure 1.
3.3 Interpolation on tree-based FTN
The goal of this subsection is to prepare a target function $M(c)$ in the FHT-W ansatz. Assuming this is done, the observable estimation task of estimating $\mathbb{E}_{C\sim P(c,t)}\left[M(C)\right]$ can be performed by an efficient tensor diagram contraction between $P(c,t)$ and $M(c)$.
For future reference, we introduce the interpolation scheme for tree-based functional tensor networks. In this procedure, the input is the target function $M$, which has an analytic expression, and the output is an FHT or FHT-W ansatz that approximates $M$. As both the FHT and the FHT-W ansatz might be used for the Dean-Kawasaki model, we introduce a general interpolation subroutine under the general tree structure.
Example of observable function
We give an example of an observable function that one compresses in practice. Suppose one has used the FHT-W model to approximate $P(c,t)$. An important observable estimation task is based on the observable function
$M(\pi)=-\sum_{j=1}^{d}\frac{\pi_{j}}{d}\log(\frac{\pi_{j}}{d}),$
which is the entropy of $\pi$. Performing observable estimation over $M(\pi)$ is important as it measures the disorder in the system. The entropy term also appears naturally in the variational form of the Fokker-Planck equation [24] and the Dean-Kawasaki model [23]. Let $M(c)$ be the pushforward of $M(\pi)$ with $M(c)=M(\pi(c))$. To be rigorous, let $s^{-1}(\cdot)$ denote the transformation from $s$ to the $\pi$ variable, and let $\mathrm{IDWT}(\cdot)$ be the inverse wavelet transformation, and then one can write $M(c)=M(s^{-1}(\mathrm{IDWT}(c))$. One can see that the function $M(c)$ has a complicated functional form and thus can only be prepared in an FHT-W through interpolation.
The derivation in [50] shows that interpolation and density estimation in FHT ansatz can be done by solving structured independent linear equations. FHT ansatz is a special case of the tree-based FTN, and we shall show that the interpolation of a function $M(x)$ in a tree-based FTN can be done similarly by solving linear equations.
Linear equation for $G_{k}$
We assume that $M(x)$ is a tree-based FTN defined over a tree graph $T=(V,E)$. Our goal is to solve for each $G_{k}$ for $k\in V$ by a tractable linear equation. We first derive the linear equation for each $G_{k}$. Let $e_{1},\ldots,e_{\mathrm{deg}(k)}$ be all edges incident to $k$, and let $e_{j}=(v_{j},k)$ for $j=1,\ldots,\mathrm{deg}(k)$. For each $v_{j}$, we construct $D_{v_{j}\to k}$ by contracting all tensor components $G_{w}$ for $w\in v_{j}\to k$. When $k\in V_{\mathrm{ext}}=[d]$, Definition 3.1 implies the following linear equation:
(14) $D(i_{1},\ldots,i_{d})=\sum_{a_{\mathcal{E}(k)}}G_{k}(i_{k},a_{\mathcal{E}(k)})\prod_{j=1}^{\mathrm{deg(k)}}D_{v_{j}\to k}(\alpha_{e_{j}},i_{v_{j}\to k\cap[d]}),$
and the equation for $k=d+1,\ldots,\tilde{d}$ can be obtained from Equation 14 by simply omitting the $i_{k}$ index in the right hand side of the equation.
Sketched equation for $G_{k}$
We see that Equation 14 is exponential-sized and is thus not practical. Overall, one solves for $G_{k}$ by solving a sketched linear equation. We first consider the case for $k\in V_{\mathrm{ext}}=[d]$. For $j=1,\ldots,\mathrm{deg}(k)$, we let $\tilde{r}_{e_{j}}>r_{e_{j}}$ be some integer, and we introduce sketch tensors $S_{v_{j}\to k}\colon\prod_{l\in v_{j}\to k\cap[d]}[n_{l}]\times[\tilde{r}_{e_{j}}]$. By contracting Equation 14 with $\bigotimes_{j=1}^{\mathrm{deg}(k)}S_{v_{j}\to k}$, one obtains the following sketched linear equation for $G_{k}$:
(15) $B_{k}(i_{k},\beta_{\mathcal{E}(k)})=\sum_{\alpha_{\mathcal{E}(k)}}\prod_{j=1}^{\mathrm{deg}(k)}A_{v_{j}\to k}(\beta_{e_{j}},\alpha_{e_{j}})G_{k}(i_{k},\alpha_{\mathcal{E}(k)}),$
where $B_{k}$ is the contraction of $D$ with $\bigotimes_{j=1}^{\mathrm{deg}(k)}S_{v_{j}\to k}$, and each $A_{v_{j}\to k}$ is the contraction of $D_{v_{j}\to k}$ with $S_{v_{j}\to k}$. The equation for $k=d+1,\ldots,\tilde{d}$ can likewise be obtained from Equation 15 by simply omitting the $i_{k}$ index on both sides.
The choice of sketch tensor
It suffices to specify the sketch tensors $S_{v\to v^{\prime}}$. We assume for concreteness that $f\colon[-1,1]^{d}\to\mathbb{R}$. For each edge $(v,k)$, we select $(v\to k\cap[d])$-indexed points $\{w^{\beta}_{v\to k\cap[d]}=(w^{\beta}_{j})_{j\in v\to k\cap[d]}\}_{\beta\in[\tilde{r}_{(v,k)}]}\subset[-1,1]^{|v\to k|\cap[d]}$. Corresponding to the selected points, we construct sketch tensor $S_{v\to k}\colon\prod_{j\in v\to k\cap[d]}[n_{j}]\times[\tilde{r}_{(v,k)}]\to\mathbb{R}$ to be defined by
(16) $S_{v\to k}(:,\beta)=\bigotimes_{j\in v\to k\cap[d]}\vec{\Psi}_{j}\left(w_{j}^{\beta}\right).$
Moreover, for each node $k=1,\ldots,d$, we select points $\{w_{k}^{\mu_{k}}\}_{\mu_{k}\in[\tilde{n}_{k}]}\subset[-1,1]$ for $\tilde{n}_{k}>n_{k}$. Corresponding to the selected points, we construct the sketch tensor $S_{k}\colon[\tilde{n}_{k}]\times[n_{k}]\to\mathbb{R}$ to be defined by
(17) $S_{k}(\mu_{k},i_{k})=\psi_{i_{k};k}(w_{k}^{\mu_{k}}).$
Obtain $B_{k}$ by $M$ queries
The FTN evaluation of $M$ defined in terms of $D$ is written as follows:
(18) $M(x_{1},\ldots,x_{d})=\sum_{i_{k}\in[n_{k}],k\in[d]}D_{i_{1},\ldots,i_{d}}\psi_{i_{1},1}(x_{1})\cdots\psi_{i_{d},d}(x_{d})=\left<D,\,\bigotimes_{j=1}^{d}\vec{\Psi}_{j}(x_{j})\right>.$
To obtain $B_{k}$, one contracts $C$ with $\otimes_{j=1}^{\mathrm{deg}(k)}S_{v_{j}\to k}$. Write $\beta_{j}=\beta_{(w_{j},k)}$ for $j=1,\ldots,\mathrm{deg}(k)$. When $k=1,\ldots,d$, combining Equation 16 and Equation 18 implies
(19) $\sum_{i_{k}}S_{k}(\mu_{k},i_{k})(w_{k}^{\mu_{k}})B_{k}(i_{k},\beta_{\mathcal{E}(k)})=M(w^{\beta_{1}}_{v_{1}\to k\cap[d]},\ldots,w^{\beta_{\mathrm{deg}(k)}}_{v_{\mathrm{deg}(k)}\to k\cap[d]},w^{\mu_{k}}_{k}),$
and so one can obtain $B_{k}$ as the unique least-squares solution to Equation 19. In practice, we directly use the pseudo-inverse formula
(20) $B_{k}(i_{k},\beta_{\mathcal{E}(k)})=\sum_{\mu_{k}}S_{k}^{\dagger}(i_{k},\mu_{k})M(w^{\beta_{1}}_{(v_{1}\to k)\cap[d]},\ldots,w^{\beta_{\mathrm{deg}(k)}}_{(v_{\mathrm{deg}(k)}\to k)\cap[d]},w^{\mu_{k}}_{k}).$
When $k=d+1,\ldots,\tilde{d}$, directly using Equation 16 and Equation 18 implies
(21) $B_{k}(\beta_{\mathcal{E}(k)})=M(w^{\beta_{1}}_{(v_{1}\to k)\cap[d]},\ldots,w^{\beta_{\mathrm{deg}(k)}}_{(v_{\mathrm{deg}(k)}\to k)\cap[d]}).$
Obtain $A_{k}$ by $M$ queries + SVD
To obtain $A_{(k,v)}$, we take the approach in [54] whereby we first obtain $Z_{(k,v)}$ from contracting $C$ with $S_{k\to v}\otimes S_{v\to k}$. The contraction can be done by querying $M$, as Equation 18 and Equation 16 implies the equation
(22) $Z_{(k,v)}(\beta,\gamma)=M(w^{\beta}_{k\to v},w^{\gamma}_{v\to k}).$
One then obtains $A_{k\to v},A_{v\to k}$ through SVD of $Z_{(k,v)}$, resulting in $Z_{(k,v)}(\beta,\gamma)=\sum_{\mu}U(\beta,\mu)W(\mu,\gamma)$. The choice of $A_{k\to v}=U$ and $A_{v\to k}=W$ forms a unique and consistent choice of gauge between the pair $(D_{v\to k},D_{k\to v})$. As obtaining $U,W$ requires merging the middle factor of the SVD to either the left or right factor, one can use the root information to specify $U,W$ fully. In particular, if $v$ is the parent to $k$, then one takes $U$ to be the left orthogonal factor in the SVD. If $k$ is the parent to $v$, then one takes $W$ to be the right orthogonal factor in the SVD. Finally, one obtains $A_{k}$ by tensor products of all $A_{v\to k}$ for $v\in\mathcal{N}(k)$.
Algorithmic summary
We summarize the tree-based FTN interpolation algorithm in Algorithm 1.
Algorithm 1 Tree-based FTN interpolation.
0:  Access to $M\colon[-1,1]^{d}\to\mathbb{R}$.
0:  Tree $T=(V,E)$ with $V=[\tilde{d}]$ and $V_{\mathrm{ext}}=[d]$.
0:  Chosen variable-dependent function basis $\{\vec{\Psi}_{j}\}_{j\in[d]}$.
0:  Collection of sketch tensors $\{S_{v\to k},S_{k\to v}\}$ and target internal ranks $\{r_{(k,v)}\}$ for each edge $(k,v)\in E$.
1:  for each edge $(k,v)$ in $T$ do
2:     Obtain $Z_{(k,v)}$ by Equation 22.
3:     Obtain $A_{k\to v}$ as the left factor of the best rank $r_{(k,v)}$ factorization of $Z_{(k,v)}$
4:     Obtain $A_{v\to k}$ as the right factor of the best rank $r_{(k,v)}$ factorization of $Z_{(k,v)}$.
5:  end for
6:  for each node $k$ in $T$ do
7:     if node $k$ contains a physical bond then
8:        Obtain $B_{k}$ by Equation 20.
9:     else
10:        Obtain $B_{k}$ by Equation 21.
11:     end if
12:     With $\mathcal{E}(k)=\{(v_{j},k)\}_{j=1}^{\mathrm{deg}(k)}$, collect each $A_{v_{j}\to k}$.
13:     Obtain $G_{k}$ by solving the over-determined linear system $(\bigotimes_{j=1}^{\mathrm{deg}(k)}A_{v_{j}\to k})G_{k}=B_{k}$.
14:  end for
4 Numerical experiment
We present the numerical performance of the proposed method for the 1D and 2D Dean-Kawasaki models. Section 4.1 covers the result for 1D. Section 4.2 covers the result for 2D.
4.1 1D Dean-Kawasaki models
Dean-Kawasaki without potential function
We first consider a simple model by taking $V_{1}=V_{2}=0$ in Equation 1. In this case, there is no external field potential or pairwise interaction. This simple case has been of research interest and has been studied in [7]. The model writes
$\partial_{t}\pi=\frac{1}{\beta}\Delta\pi+\mathrm{div}(\sqrt{\frac{2\pi}{\beta N}}\eta).$
We take $\beta=0.05$, $h=\frac{1}{64}$, and $N=1000$. The samples are generated by sampling the SDE according to Equation 7 with $\delta t=0.005$. In this case, the dimension is $d=64$. We are interested in the solution of the Fokker-Planck equation at $t=1$, and we take the initial condition to be $\pi(x,0)=1$ for all $x\in[0,1]$. We collect $B=6000$ samples with the procedure in Section 2.1 and perform a transformation to the wavelet coefficient $c$ as described in Section 3.1. Based on the samples, we use the density estimation procedure as given in [54] for the FHT-W model. We take the maximal internal dimension to be $r=20$. On each variable, we use a Legendre polynomial basis with a maximal degree of $q=25$.
Figure 3: 1D Dean-Kawasaki model with $V_{1}=V_{2}=0$. Plot of the correlation matrix predicted by the FHT-W ansatz.
Figure 4: 1D Dean-Kawasaki model with an external potential $V_{1}$ and a pairwise potential $V_{2}$ modeling particle repulsion. Plot of the correlation matrix predicted by the FHT-W ansatz.
We first use approximated density in the FHT-W ansatz to predict the entropy
$M(\pi)=-\sum_{j=1}^{d}\frac{\pi_{j}}{d}\log(\frac{\pi_{j}}{d}).$
As mentioned in Section 3.3, to perform observable estimation by tensor contraction, we need to construct the function $M(c)=M(\pi(c))$. We use Algorithm 1 on $M(c)$ for interpolation. On each variable, we use a Legendre polynomial basis with a maximal degree of $q=6$, and we use a maximal rank of $r=5$. The result of FHT-W interpolation is in Figure 5, which shows that the interpolation of $M(c)$ is highly accurate.
Figure 5: Interpolations for 1D Dean-Kawasaki model and 2D Dean-Kawasaki model without potential functions. The target function is $M(c)=M(s^{-1}(\mathrm{IDWT}(c))$, where $M(\pi)$ is the Shannon entropy function. The plot compares the function evaluation between $M(c)$ and the FHT-W ansatz obtained from the interpolation procedure in Algorithm 1. The $20000$ evaluation points are chosen randomly. The mean relative prediction error is $8.7\times 10^{-9}$ for 1D Dean-Kawasaki and $1.7\times 10^{-6}$ for 2D Dean-Kawasaki.
The predicted entropy obtained after interpolation is $4.15827$, while the true reference entropy from the Monte Carlo estimation is $4.15825$. The relative FHT-W prediction error on $M(\pi)$ is $5.3\times 10^{-6}$. We repeat the experiments for the 2-Rényi entropy $M(\pi)=-\log(\sum_{j=1}^{d}\pi_{j}^{2}/d^{2}),$ and we likewise obtain a relative prediction error of $1.0\times 10^{-5}$. For both observables, using the obtained approximated density for observable estimation is on the same order of accuracy as directly estimating $M(\pi)$ through the $B=6000$ samples.
To give a more challenging numerical test, we use the FHT-W ansatz to predict the correlation matrix,
$R(i,i^{\prime})=\mathrm{Corr}_{\Pi\sim P(\pi,t)}\left(\Pi_{i},\Pi_{i^{\prime}}\right),$
which can be carried out by performing $O(d^{2})$ observable estimation tasks. Figure 4 shows that the predicted correlation closely matches the ground truth obtained from sample estimation.
Dean-Kawasaki with external potential and repulsion
We consider the Dean-Kawasaki model with an external potential term $V_{1}$ and a pairwise potential term $V_{2}$. In this case, the equation writes
$\partial_{t}\pi=\frac{1}{\beta}\Delta\pi+\mathrm{div}\left(\sqrt{\frac{2\pi}{\beta N}}\eta+\pi\nabla V_{1}+\pi\nabla V_{2}*\pi\right),$
where take $\beta=0.05$, $h=\frac{1}{64}$, $N=1000$. For the external potential, we take
$V_{1}(x)=-15\cos(2\pi(x-0.5)),$
which leads the distribution $\pi(x,t)$ to be more concentrated around $x\approx 0.5$.
For the pairwise potential, we take $V_{2}$ to be a regularized repulsion term written as follows:
$V_{2}(x)=-\frac{6}{x^{2}+0.01},$
which is a smooth penalization term. Due to the periodic boundary condition, for two locations $z,z^{\prime}\in[0,1]$, we choose their relative distance to be $x=\min(|z-z^{\prime}|,1-|z-z^{\prime}|)$, and $V_{2}(x)$ is the term for the two-body interaction term. We remark that $V_{2}(x)$ and $V_{2}^{\prime}(x)$ is insignificant unless $x$ is close to 00, and so the boundary effect from $x$ is negligible.
In this case, the dimension is $d=64$. The samples are generated by sampling the SDE according to Equation 7 with $\delta t=0.0002$. We are interested in the Fokker-Planck equation at $t=1$, and we take the initial condition to be $\pi(x,0)=1$ for all $x\in[0,1]$. We collect $B=6000$ samples and perform a transformation to the $c$ coordinate as described in Section 3.1. Based on the samples, we use the density estimation procedure for the FHT-W model with the maximal internal dimension taken to be $r=20$. On each variable, we use a Legendre polynomial basis with a maximal degree of $q=25$.
For this case, we repeat the procedure for estimating the Shannon entropy and the 2-Rényi entropy. In this case, compressing $M(\pi)$ in the $c$ coordinate as an FHT-W ansatz is also highly accurate. For the Shannon entropy, the relative prediction error is $1.3\times 10^{-5}$. For the 2-Rényi entropy, the relative prediction error is $1.4\times 10^{-5}$. In both cases, using the obtained approximated density for observable estimation is on the same order of accuracy as directly estimating $M(\pi)$ through the $B=6000$ samples.
Finally, we use the FHT-W model to predict the correlation matrix. Figure 4 shows that the FHT-W ansatz quite accurately predicts the correlation matrix $R$.
4.2 2D Dean-Kawasaki models
Dean-Kawasaki without potential function
We first consider the Dean-Kawasaki model with $V_{1}=V_{2}=0$
$\partial_{t}\pi=\frac{1}{\beta}\Delta\pi+\mathrm{div}(\sqrt{\frac{2\pi}{\beta N}}\eta),$
where we take $\beta=0.05$, $h=\frac{1}{8}$, and $N=1000$. In this case, the dimension is $d=64$. We are interested in the solution of the Fokker-Planck equation at $t=1$, and we take the initial condition to be $\pi(x,0)=1$ for all $x\in[0,1]^{2}$. We collect $B=12000$ samples by performing a time discretization of the SDE in Section 2.2, and in particular, we take $\delta t=0.001$. We then perform a transformation to the $c$ coordinate as described in Section 3.1. Based on the samples, we use the density estimation procedure as given in [54] for the FHT-W model. We take the maximal internal dimension to be $r=20$. On each variable, we use a Legendre polynomial basis with a maximal degree of $q=15$.
We first test the observable estimation in terms of the Shannon entropy and the 2-Rényi entropy. The 2D wavelet transform leads to a more complicated functional form for $M(c)=M(s^{-1}(\mathrm{IDWT}(c))$. We show the result of FHT-W interpolation in Figure 5, which shows that the interpolation of $M(c)$ is highly accurate. For the Shannon entropy, the relative prediction error is $5.5\times 10^{-6}$. For the 2-Rényi entropy, the relative prediction error is $4.5\times 10^{-5}$.
Lastly, we use the obtained FHT-W ansatz to predict the value of the two-point correlation function,
$f(i,j)=\mathrm{Corr}_{\Pi\sim P(\pi,t)}\left(\Pi_{(i,j)},\Pi_{(4,4)}\right),$
which is done by using the obtained ansatz to perform repeated observable estimation tasks. Figure 7 shows that the proposed FHT-W ansatz closely matches the ground truth.
Figure 6: 2D Dean-Kawasaki model with $V_{1}=V_{2}=0$. The plot of the two-point correlation function predicted by the FHT-W ansatz.
Figure 7: 2D Dean-Kawasaki model with an external potential $V_{1}$ and a pairwise potential term $V_{2}$ modeling particle repulsion. The plot of the two-point correlation function predicted by the FHT-W ansatz.
Dean-Kawasaki with external potential and repulsion
We consider the 2D Dean-Kawasaki model with an external potential term $V_{1}$ and a pairwise potential term $V_{2}$:
$\partial_{t}\pi=\frac{1}{\beta}\Delta\pi+\mathrm{div}\left(\sqrt{\frac{2\pi}{\beta N}}\eta+\pi\nabla V_{1}+\pi\nabla V_{2}*\pi\right),$
where take $\beta=0.05$, $h=\frac{1}{8}$, $N=1000$. For the external potential and pairwise potential, we take
$V_{1}(x_{1},x_{2})=-15\cos(2\pi(x_{1}-0.5))-15\cos(2\pi(x_{2}-0.5)),$
which leads the distribution $\pi(x,t)$ to be more concentrated around $x\approx(0.5,0.5)$. For the pairwise interaction, we take
$V_{2}(x_{1},x_{2})=-\frac{800}{x_{1}^{2}+0.1}-\frac{800}{x_{2}^{2}+0.1},$
which leads to substantial repulsion for the resulting interacting particle scheme. In this case, the dimension is $d=64$. We are interested in the solution of the Fokker-Planck equation at $t=1$, and we take the initial condition to be $\pi(x,0)=1$ for all $x\in[0,1]^{2}$. The samples are generated by sampling the SDE according to Equation 9 with $\delta t=0.0003$.
For the density estimation, we repeat the setting for the previous 2D case with $B=12000$, maximal bond dimension $r=20$. On each variable, we use a Legendre polynomial basis with a maximal degree of $q=15$. For the Shannon entropy and 2-Rényi entropy, we respectively obtain a relative error of $3.5\times 10^{-5}$ and $1.3\times 10^{-4}$.
Lastly, we use the obtained FHT-W ansatz to predict the value of the two-point correlation function. Figure 7 shows that the proposed FHT-W ansatz closely matches the ground truth.